OpenStudy (peterg2001):

A quantity, y, varies directly as x. When y = 10, x = 6. Find x when y = 14. 4.3 8.4 23.3 60

OpenStudy (anonymous):

poo

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

ok so x=6 and y=10 and you need to find it

OpenStudy (anonymous):

Find x when y=14 x=6 y=10 Find 6 when 10=14

OpenStudy (godlovesme):

Y varies directly as x means y=x .oh nmv @king_trade_OKC will help u :)

OpenStudy (anonymous):

23.3

OpenStudy (anonymous):

nah thats as far as a could go... You got it @Godlovesme

OpenStudy (radar):

Step 1. Determine the "constant of proprotionality" k. We are given the y varies directly with x or y=kx We are further given that 10 = k6 or k=10/6 Step 2, Determine x from the 2nd relationship or 14 = 10x/6 just solve for x.

OpenStudy (anonymous):

all you have to do really is find the other x, if it gives you 2 y varies, there's not much else to but cross multiply and divide, isn't there? giving you 10 x 14 = 140/6 = 23.33333 so forth.

OpenStudy (radar):

The solution is actually 8.4